What is the quotient of 2 4/5 divided by 1 1/3
Answer by tiffany222(56)

2 4/5 divided by 1 1/3
= 14/5 divided by 4/3
= 14/5 * 3/4
= 42/20
= 21/10
= 2 1/10
Final Answer: 2 1/10